OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 2 and a y-intercept of -3.

OpenStudy (raden):

use this : y = mx + c with m = slope and c = y-intercept

OpenStudy (anonymous):

i thought it was mx+b?

OpenStudy (raden):

though one :)

OpenStudy (anonymous):

ikr... i really cant do it :(

OpenStudy (raden):

y = mx + b here given m=2 and b=-3

OpenStudy (anonymous):

so what is X?

OpenStudy (raden):

it's just a variable

OpenStudy (anonymous):

oh k can u just solve it for me koz i c black on black

OpenStudy (raden):

y = mx + b, u just put the values of m=2 and b=-3, it can be y = 2x - 3

OpenStudy (anonymous):

thats the result right?

OpenStudy (raden):

yep

OpenStudy (anonymous):

u deserve a medal

OpenStudy (raden):

thank you ;)

OpenStudy (anonymous):

welc :)